OpenStudy (anonymous):

Paul biked 14.6 miles on Saturday and 12.8 miles on Sunday. Mariah biked the same total distance as Paul, but only biked 11.7 miles on Saturday. How far did Mariah bike on Sunday? A. 27.8 miles B. 1.8 miles C. 15.7 miles D. 13.5 miles

OpenStudy (anonymous):

14.6 + 12.8 = 27.4 (total miles biked by Paul and Mariah) x - 11.7 = 27.4 (Solve for x which is the total miles biked by Mariah in Sunday.)

OpenStudy (anonymous):

Wut

OpenStudy (anonymous):

AH

OpenStudy (anonymous):

*x + 11.7 = 27.4

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

I have a few more if you don't mind @Sadworld

OpenStudy (anonymous):

I mean it's addition. Typo

OpenStudy (anonymous):

But c is correct

OpenStudy (anonymous):

Ok :P

OpenStudy (anonymous):

Just tag me to your questions. I'll help

OpenStudy (anonymous):

A rectangle has a length of 10 inches and a perimeter of 48 inches. Which equation could be used to find the width of the rectangle? A. 48 = 20 + 2w B. 48 = 20 – 2w C. 48 = 40w D. 48 = 10w @Sadworld

OpenStudy (anonymous):

Also @Sadworld If you have skype it would help alot since i have a whole 1-2 units to do

OpenStudy (anonymous):

I got skype, just give a moment to dig for it.

OpenStudy (anonymous):

Kk

OpenStudy (anonymous):

Pm me your skype user

OpenStudy (anonymous):

Okay, it's deeper than I thought. Got to dig. Give me few more mins. Perimeter = 2(length)+2(width) 48 = 2(10) + 2(w) 48 = 20+2w

OpenStudy (anonymous):

Ok :D

OpenStudy (anonymous):

\[14.6\, +12.8\, -11.7=15.7\]

OpenStudy (anonymous):

@robtobey Answered already :)

OpenStudy (anonymous):

Thanks anyways @robtobey

OpenStudy (anonymous):

Your Welcome.